# Mitigating Hallucinations in Vision-Language Models through Image-Guided Head Suppression

### v1

@@ A.4 Extended Evaluation of SPIN
Model | Decoding | Method | Layers | Supp. | Scale | CHAIRS | CHAIRI | F1 |
Heads | Factor | () | () | () | ||||
Minigpt4 | Beam Search | Baseline | - | - | - | 33.4 | 11.0 | 70.2 |
OPERA | - | - | - | 23.2 | 8.8 | 69.4 | ||
PAI | - | - | - | 22.2 | 8.3 | 68.1 | ||
SPIN | 124 | 0.15 | 0.0 | 16.6 | 4.8 | 68.2 | ||
SPIN | 116 | 0.18 | 0.05 | 16.2 | 6.6 | 69.1 | ||
Nucleus Sampling | Baseline | - | - | - | 31.8 | 11.8 | 66.4 | |
PAI | - | - | - | 23.0 | 9.0 | 68.3 | ||
SPIN | 116 | 0.18 | 0.05 | 23.6 | 12.5 | 62.9 | ||
Shikra | Beam Search | Baseline | - | - | - | 52.6 | 14.1 | 75.7 |
OPERA | - | - | - | 40.2 | 12.6 | 72.7 | ||
PAI | - | - | - | 38.6 | 10.1 | 74.9 | ||
SPIN | 132 | 0.40 | 0.0 | 32.4 | 10.0 | 74.5 | ||
Nucleus Sampling | Baseline | - | - | - | 56.0 | 15.7 | 72.8 | |
VCD | - | - | - | 55.6 | 15.2 | 74.9 | ||
PAI | - | - | - | 33.8 | 8.7 | 75.1 | ||
SPIN | 132 | 0.40 | 0.0 | 34.2 | 10.6 | 71.5 |
@@
POPE Evaluation: We assess SPIN on POPE evaluation to examine its performance in multi-turn dialogue settings. All experiments in this section are performed using greedy decoding. As shown in Table 8, SPIN consistently reduces hallucinations across multiple models. These results demonstrate that our method not only generalizes well beyond CHAIR, but also achieves strong hallucination mitigation on multi-turn vision-language tasks.
@@
CHAIR Results under Diverse Decoding Strategies: To further evaluate the effectiveness of SPIN, we conduct experiments on CHAIR under various decoding strategies: greedy search, beam search, and nucleus sampling. Across four models: LLaVA-1.5 (7B), MiniGPT4, Shikra, and Qwen-VL. We explore a wider range of configurations by varying the ratio of suppressed heads, scaling factors, and the layers where SPIN is applied.
@@
As discussed in Section 4.2, we have already provided a detailed analysis of SPIN’s performance under greedy search for all four models, as well as its behavior on beam search and nucleus sampling for LLaVA-1.5 (7B) and Qwen-VL. Here, we highlight the results for MiniGPT4 and Shikra under beam search and nucleus sampling, as shown in Table 9. For MiniGPT4, SPIN reduces CHAIR and CHAIR by 2.3 and 2.0 with beam search. For Shikra, the corresponding reductions are 1.4 and 1.6. Under nucleus sampling, SPIN yields less significant improvements, but still effectively mitigates hallucinations. For MiniGPT4, SPIN reduces CHAIR by 1.35. On Shikra, it reduces CHAIR and CHAIR by 1.5 and 1.6, with only a 1.3% drop in F1.
A.5 POPE Visualization
In Figure 8, we visualize outputs for POPE evaluation on random and adversarial splits. SPIN successfully mitigates hallucinations in both cases.
@@ A.6 CHAIR Visualization
The captions generated using the baseline and SPIN are visualized in Figures 9, 10, 11, 12, and 13, for LLaVA-1.5 (7B), LLaVA-1.5 (13B), Shikra, Qwen-VL, and Minigpt4. SPIN effectively mitigates hallucinations generated by the baseline models.
GPT-4o Prompt |
You are required to score the performance of two AI assistants in describing a given image. You should pay extra attention to the hallucination, which refers to the part of descriptions that are inconsistent with the image content, such as claiming the existence of something not present in the image or describing incorrectly in terms of the counts, positions, or colors of objects in the image. Please rate the responses of the assistants on a scale of 1 to 10, where a higher score indicates better performance, according to the following criteria: 1: Accuracy: whether the response is accurate with respect to the image content. Responses with fewer hallucinations should be given higher scores. 2: Detailedness: whether the response is rich in necessary details. Note that hallucinated descriptions should not count as necessary details. Please output the scores for each criterion, containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. Following the scores, please provide an explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment. [Assistant 1] {Response of Assistant 1} [End of Assistant 1] [Assistant 2] {Response of Assistant 2} [End of Assistant 2] Output format: Accuracy: <Scores of the two answers> |
Reason: Detailedness: <Scores of the two answers> |
Reason: |

### v2

@@ A.4 Extended Evaluation of SPIN
Model
Decoding
Method
Layers
Supp.
Scale
CHAIRS
CHAIRI
F1
Heads
Factor
MiniGPT-4
Beam
Search
Baseline
33.4
11.0
70.2
OPERA
23.2
8.8
69.4
PAI
22.2
8.3
68.1
SPIN
124
0.15
0.0
16.6
4.8
68.2
SPIN
116
0.18
0.05
16.2
6.6
69.1
Nucleus
Sampling
Baseline
31.8
11.8
66.4
PAI
23.0
9.0
68.3
SPIN
116
0.18
0.05
23.6
12.5
62.9
Shikra
Beam
Search
Baseline
52.6
14.1
75.7
OPERA
40.2
12.6
72.7
PAI
38.6
10.1
74.9
SPIN
132
0.40
0.0
32.4
10.0
74.5
Nucleus
Sampling
Baseline
56.0
15.7
72.8
VCD
55.6
15.2
74.9
PAI
33.8
8.7
75.1
SPIN
132
0.40
0.0
34.2
10.6
71.5
Table 11: CHAIR evaluation using MiniGPT-4 and Shikra under beam search and nucleus sampling decoding.
POPE Evaluation:
@@
We assess SPIN on POPE evaluation to examine its performance in multi-turn dialogue settings. All experiments in this section are performed using greedy decoding. As shown in Table 10, SPIN consistently reduces hallucinations across multiple models. These results demonstrate that our method not only generalizes well beyond CHAIR, but also achieves strong hallucination mitigation on multi-turn vision-language tasks.
CHAIR Results under Diverse Decoding Strategies:
@@
To further evaluate the effectiveness of SPIN, we conduct experiments on CHAIR under various decoding strategies: greedy search, beam search, and nucleus sampling. Across four models: LLaVA-1.5 (7B), MiniGPT-4, Shikra, and Qwen-VL. We explore a wider range of configurations by varying the ratio of suppressed heads, scaling factors, and the layers where SPIN is applied.
@@
As discussed in Section 4.2, we have already provided a detailed analysis of SPIN’s performance under greedy search for all four models, as well as its behavior on beam search and nucleus sampling for LLaVA-1.5 (7B) and Qwen-VL. Here, we highlight the results for MiniGPT-4 and Shikra under beam search and nucleus sampling, as shown in Table 11. For MiniGPT-4, SPIN reduces CHAIR and CHAIR by 2.3 and 2.0 with beam search. For Shikra, the corresponding reductions are 1.4 and 1.6. Under nucleus sampling, SPIN yields less significant improvements, but still effectively mitigates hallucinations. For MiniGPT-4, SPIN reduces CHAIR by 1.35. On Shikra, it reduces CHAIR and CHAIR by 1.5 and 1.6, with only a 1.3% drop in F1.
Model
Method
Layers
CS()
CI()
F1()
LLaVA-
Next (7B)
Baseline
35.6
8.6
71.7
SPIN
132
0.05
0.0
32.6
8.1
72.3
SPIN
116
0.20
0.05
26.8
7.4
71.7
Qwen2.5-
VL (7B)
Baseline
34.8
7.9
75.7
SPIN
116
0.02
0.0
28.6
7.0
74.0
SPIN
116
0.02
0.05
30.6
7.5
74.7
Table 12: CHAIR evaluation results on the latest LVLMs, including LLaVA-Next and Qwen2.5-VL.
Model
Mode
Layers
Accuracy
F1
LLaVA-Next
Baseline
88.45
87.99
SPIN
132
0.05
0.0
89.00
88.52
SPIN
132
0.05
0.001
89.07
88.60
Qwen2.5-VL
Baseline
81.64
77.75
SPIN
120
0.02
0.0
83.02
79.80
SPIN
120
0.02
0.001
83.05
79.87
Table 13: Experimental results on Multi-turn POPE using LLaVA-Next and Qwen2.5-VL.
A.5 Evaluations on the Latest Models
To further assess the accessibility and robustness of our method on state-of-the-art LVLMs, we conduct evaluations under greedy decoding on LLaVA-Next Liu et al. (2024c) and Qwen2.5-VL Bai et al. (2025). The CHAIR results are reported in Table 12, while the POPE results are shown in Table 13. SPIN consistently alleviates hallucinations in these newer models and provides improved accuracy and F1 for POPE evaluation.
It is worth noting that, due to differences in architectures in the language backbone, Qwen2.5-VL generally benefits more from suppressing a smaller fraction of heads. Specifically, Qwen2.5-VL comprises 28 layers with 28 heads per layer, whereas LLaVA-Next, leveraging Mistral-7B, consists of 32 layers with 32 heads each. This indicates that each attention head in Qwen2.5-VL undertakes greater functional responsibility, so overly aggressive suppression of heads can undermine the model’s capabilities.
@@ A.6 CHAIR Visualization
The captions generated using the baseline and SPIN are visualized in Figures 8, 9, 10, 11, and 12, for LLaVA-1.5 (7B), LLaVA-1.5 (13B), Shikra, Qwen-VL, and MiniGPT-4. SPIN effectively mitigates hallucinations generated by the baseline models.
Table 14: The prompt used for GPT-4o evaluation adopted from Liu et al. (2024d).
GPT-4o Prompt
You are required to score the performance of two AI assistants in describing a given image. You should pay extra attention to the hallucination, which refers to the part of descriptions that are inconsistent with the image content, such as claiming the existence of something not present in the image or describing incorrectly in terms of the counts, positions, or colors of objects in the image. Please rate the responses of the assistants on a scale of 1 to 10, where a higher score indicates better performance, according to the following criteria:1: Accuracy: whether the response is accurate with respect to the image content. Responses with fewer hallucinations should be given higher scores.2: Detailedness: whether the response is rich in necessary details. Note that hallucinated descriptions should not count as necessary details.Please output the scores for each criterion, containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. Following the scores, please provide an explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.[Assistant 1]{Response of Assistant 1}[End of Assistant 1][Assistant 2]{Response of Assistant 2}[End of Assistant 2]Output format:Accuracy: <Scores of the two answers>
Reason:Detailedness: <Scores of the two answers>
Reason:
Figure 8: SPIN’s effectiveness in reducing hallucinations with LLaVA-1.5 (7B).Figure 9: SPIN’s effectiveness in reducing hallucinations with LLaVA-1.5 (13B).Figure 10: SPIN’s effectiveness in reducing hallucinations with Shikra.Figure 11: SPIN’s effectiveness in reducing hallucinations with Qwen-VL-Chat.Figure 12: SPIN’s effectiveness in reducing hallucinations with MiniGPT-4.